# UltraClean: A Simple Framework to Train Robust Neural Networks against Backdoor Attacks

### v1

@@ 1 Introduction
With the thriving of machine learning, deep neural networks (DNNs) have achieved unprecedented progress and reached human-level performance in various tasks, including computer vision [31, 18], natural language processing [14, 5] and game playing [58]. Given the remarkable success, DNNs are further deployed to safety-critical applications such as authentication [65] and autonomous driving [4]. However, DNNs are proven to be vulnerable to a variety type of adversarial attacks. One notorious attack is the adversarial example [27] that occurs at test-time, where an adversary fools well-trained DNN models by adding imperceptible perturbations on test images. Another well-known attack is the data poisoning attack [3] that occurs in the training phase, where an adversary can inject well-crafted poisoned samples into the training data and introduce malicious behavior to models trained on the poisoned dataset. This paper focuses on backdoor attacks [28], where an adversary attempts to contaminate the training dataset via data poisoning and install a backdoor into models trained on the corrupted dataset. During inference, backdoored models misclassify inputs with backdoor triggers to a target label while behaving normally on benign inputs.
@@
Data poisoning backdoor attacks come in two flavors: dirty-label attacks [10, 49, 51] and clean-label attacks [2, 62, 57]. Poisoned samples of dirty-label attacks are always mislabeled while clean-label attacks hold consistent labels to images content. Although dirty-label attacks are effective, they can be easily distinguished due to incorrect labels. Existing defenses have shown decent performance in detecting and mitigating such attacks [26, 43]. On the other hand, clean-label attacks are more stealthy and insidious. The clean-label poisoned samples are almost visually indistinguishable from benign samples; thus, defenses against the attack become a more challenging task. Recent works [41, 33] attempted to alleviate the clean-label attacks by decoupling the training phase to suppress the backdoor injection. However, they require complicated algorithms with significantly more operations in training and do not identify the backdoor samples in the poisoned dataset. Thus, users have to re-run the defense algorithms every time they train a new model by using the same potentially poisoned dataset, which dramatically increases the cost.
@@
In this work, we propose UltraClean, a poisons-filtering-based framework to detect backdoor samples, cleanse poisoned datasets, and train backdoor-mitigated models against both dirty-label and clean-label backdoor attacks. We focus on the image classification task. In contrast to the aforementioned poisons-suppression-based defenses, poisons-filtering-based defenses only require a one-time cleanse of the dataset. Users need to run the proposed algorithm only once to obtain the sanitized dataset, and any model trained on the sanitized dataset would be backdoor-free. Our idea is inspired by prior works [44] and [67, 68], which have demonstrated that adversarial perturbations of adversarial examples are amplified during the feed-forward propagation (i.e., error amplification effect) in DNNs and can be effectively eliminated by simple image-denoising techniques. We argue that although backdoor samples hold fundamentally different generation mechanisms to adversarial examples, the backdoor triggers share a similar error amplification effect as the adversarial perturbations. To this end, our training framework, UltraClean, employs off-the-shelf image-denoising functions and the error amplification effect to filter out poisoned samples from benign training data and thwart the backdoor implantation. The proposed method first trains an arbitrary model on the potentially poisoned dataset and uses it as the backdoor detection model. We then produce two variants of each training image using denoising functions and feed the difference corresponding to the original image into the detection model. We leverage the error amplification effect to compute the susceptibility of the training data in a feed-forward pass. The susceptibility of poisoned data tends to be higher, guiding the model to detect and remove these samples. Finally, we obtain a backdoor-mitigated model by retraining on the sterilized dataset. We show that UltraClean is highly effective in defending against various representative attacks including both dirty-label and clean-label ones, achieving a high detection rate while maintaining the model accuracy.
@@
  • UltraClean is a once-for-all backdoor-free training framework and a poisons-filtering-based defense against both dirty and clean-label backdoor attacks. The method not only defends against backdoor attacks but also identifies poisoned samples regardless of generation mechanisms and dataset complexity.
@@
  • We demonstrate that two state-of-the-art poisons-filtering-based defenses are only effective on dirty-label attacks but are unsuccessful in defending against clean-label attacks.
@@ 2.1 Backdoor Attacks on DNN
Backdoor attacks can be broadly categorized into data poisoning [28, 10, 57] and model poisoning [50, 51, 17, 16]. Data poisoning backdoor attacks implant the backdoors by modifying training data (i.e., injecting poisoned samples), which does not require access to the training process. Such attacks affect the victims indirectly through the poisoned dataset. In contrast, model poisoning backdoor attacks require full control of the training process to alter model parameters and embed backdoors into the model. Such attacks affect the victims by providing the poisoned model directly. Our work focuses on the defense against data poisoning attacks.
The very first backdoor attack via data poisoning against neural networks is BadNets [28] where poisoned training samples are generated by stamping a pre-defined pattern (trigger) onto benign images and assigned with incorrect target labels (dirty-label). One drawback of BadNets is that the poisoned samples are suspicious and quite easy to detect upon human inspection. Later on, a series of works are proposed to enhance the stealthiness by improving the invisibility of backdoor triggers. [10] blends backdoor triggers with benign images; [72] imposes imperceptible perturbations onto poisoned samples using universal adversarial perturbations [49]; [39, 40] creates invisible triggers via steganography and regularization; [11] employs controlled detoxification to perform the attack in the feature space. However, all these methods are still dirty-label attacks and could be detected by inspecting data labels.
On the other hand, clean-label backdoor attacks are more stealthy and hard to detect since they retain consistent labels to the image contents. For instance, [2, 47, 56] superimpose the trigger with benign images and employ different techniques (i.e., sinusoidal signal, reflection and image-scaling) to conceal the trigger; [62] creates poisoned sample by placing a stealthy patch on hard-to-classify images generated by generative models or adversarial perturbations; [57] produces poisoned samples by minimizing their distance from source benign images in feature space. The objective of UltraClean is to defend against both dirty-label attacks and clean-label attacks in a simple and effective fashion.
@@ 2.2 Defenses
Defenses against backdoor attacks are extensively studied in recent years, which can be broadly categorized as training-phase defense and inference-phase defense. Training-phase defense attempts to eliminate backdoor before model deployment by filtering poisoned training samples [7, 61, 12, 60, 24, 70, 30], suppressing the effectiveness of poisoned data [9, 23, 19, 32, 41, 33], or reconstructing the trained model [48, 45, 71, 42]. Inference-phase defense alleviates the backdoor effect after model deployment by pre-processing data [15, 63, 55], filtering suspicious samples [36, 37], synthesizing possible triggers [64, 8, 54, 29, 73], or identifying if a model is backdoored [35, 69, 34, 66, 38].
This paper proposes the first poisons-filtering-based defensive solution against a wide range of backdoor attacks (i.e., both dirty-label and clean-label). Our work is a training-phase defense where users seek to detect and remove malicious training samples and train a backdoor-free model. The most relevant works are the spectral signatures defense [61] and strong intentional perturbation defense [24] that can effectively distinguish poisoned samples in the dataset. However, we find that these defenses are only effective against dirty-label attacks.
@@ 3.1 Threat Model
We consider the same threat model defined in prior data poisoning backdoor works [61] where the adversary targets to implant a backdoor into a DNN for an image classification task by injecting a fraction of poisoned samples into the training dataset. We assume a strong adversary that knows standard training algorithms, classic model architectures, and the statistical information of the training dataset to craft powerful and stealthy poisoned samples. However, the model is not trained by the adversary but the user with possibly contaminated training dataset from an uncertified source. Under the threat model, we evaluate the effectiveness of the proposed defensive framework from three dimensions.
@@
Input: Training dataset $(\mathbf{x},\mathbf{y})$ $\sim$ $\boldsymbol{D_{tr}}$, randomly initialized deep neural network model $\mathbf{M_{\theta}}$, detection and removal threshold $\boldsymbol{{\beta}}$. Let $\boldsymbol{f_{\theta^{*}}(\cdot)}$ be the softmax output of the potentially backdoored model $\mathbf{M_{\theta^{*}}}$, $\boldsymbol{n}$ be the total number of $\boldsymbol{D_{tr}}$.
Output: Sanitized training dataset $\boldsymbol{D_{s}}$, post-clean model $\mathbf{M_{\hat{\theta}}}$
@@ 3.2 UltraClean Framework
Figure 1: Spatial and frequency domain views of two backdoor attacks. Top two rows are the dirty-label attack (BadNets [28]); Bottom two rows are the clean-label attack (Hidden Trigger Backdoor [57]). As shown in the 4th and 8th columns, poisons reveal substantial qualitative noise difference than the clean counterparts. Since it is hard to differentiate poisoned and benign samples by their difference in the pixel space, we leverage the error amplification effect of DNNs to amplify the noise difference.
@@ 3.3 Methodology
The key of UltraClean is to develop an approach that can differentiate poisoned and benign samples, which hence can be effective under both dirty-label and clean-label settings. In general, backdoor attacks require a trigger to induce DNNs to learn the malicious behavior. Such a trigger, whether visually perceptible or not, introduces extra noise (pixels that do not match the image content) to benign images. We plot the frequency-domain view in Figure 1 to visualize the signal distribution in images. The central area of the frequency-domain figure represents the low-frequency signal, while the peripheral area represents the high-frequency signal. As shown in the 1st and 5th columns, poisons demonstrate more high-frequency noise than benign data. The noise is bounded by the maximum perturbation added to the images and is sometimes hard to detect in the pixel space (i.e., some works attempt to minimize the perturbation [62, 57]). For example, it is barely impossible to perceptually distinguish the poison sample (the right side of the 3rd row) from the clean sample (the left side of the 3rd row) while they have obvious differences in the frequency domain (the 4th and 8th columns in the 4th row). Moreover, the noise distributions vary due to different mechanisms of poisoned sample generation, which demands high generalizability of the detection.
@@
Non-local mean denoising [6] removes noise by replacing each pixel value with a weighted mean computed over global spatial regions. It is defined as:
[…]
where $\tilde{c}(p)$ is the denoised value of pixel $p$, $d(\cdot)$ represents the Euclidean distance between spatial regions $\Omega(p,r)$ and $\Omega(q,r)$. $\Omega(p,r)$ and $\Omega(q,r)$ are search windows centered at pixels $p$ and $q$, and the boundary of which is defined by $r$. $C(\cdot)$ is a normalization function and $g(\cdot)$ is a decreasing function. In this work, we exploit the pixel-wise implementation of the algorithm. Consider a color image with RGB channels as $\mathbf{x}=(c_{1},c_{2},c_{3})$, the non-local mean denoising operation is expressed as:
@@
We acknowledge there exist many other denoising algorithms designed specifically for certain types of noise, which might also be used for defending against backdoor attacks. However, incorporating more denoising algorithms, especially those complicated ones, may significantly increase the complexity of our method, while providing little gain, as using the two denoising techniques can already achieve excellent performance. The workflow of UltraClean is straightforward, for each training image $\mathbf{x}$, we generate two denoised versions $\tilde{\mathbf{x_{1}}}$ (e.g., the 2nd and 6th columns in Fig 1) and $\tilde{\mathbf{x_{2}}}$ (e.g., the 3rd and 7th columns). $\tilde{\mathbf{x_{1}}}$ and $\tilde{\mathbf{x_{2}}}$ serve as baselines to compute the susceptibility. We define susceptibility as the error (i.e., noise difference) between original images and denoised variants. It can be seen from the 4th and 8th columns that the error of poisoned images are considerably higher than that of clean images. However, as seen in the 6th and 7th columns, the denoising functions do not significantly reduce the noise in the denoised variants compared to the original poisoned samples (the 5th column), so we are not able to distinguish poisoned samples from benign samples by simply computing the susceptibility in the pixel space.
We propose a method to enlarge the error to facilitate the computation using the error amplification effect, which is a property of DNNs that minor adversarial perturbations in model inputs accumulate during forward propagation and affect the model outputs [44]. We feed the original image and denoised variants into the pre-trained DNN and obtain three vectors $\mathbf{v}$, $\mathbf{v_{1}}$, and $\mathbf{v_{2}}$, which are the softmax outputs of the model. We compute the $\ell_{1}$-norm distances of ($\mathbf{v}$, $\mathbf{v_{1}}$) and ($\mathbf{v}$, $\mathbf{v_{2}}$), respectively, and then aggregate the results to obtain the amplified noise difference. The susceptibility can thus be mathematically defined as:
$\mathbf{s}=||\mathbf{v}-\mathbf{v_{1}}||_{1}+||\mathbf{v}-\mathbf{v_{2}}||_{1}$ (6)

### v2

@@ 1 Introduction
Figure 1: Illustration of dirty-label and clean-label attacks (Top row: poisoned training samples; Bottom row: backdoored test samples. Red: incorrect labels; Green: correct labels. Dirty-label poisoned samples always possess incorrect labels while clean-label poisoned samples are imperceptible compared to benign samples and possess correct labels.
@@
With the thriving of machine learning, deep neural networks (DNNs) have achieved unprecedented progress and reached human-level performance in various tasks, including computer vision [he2016deep, DBLP:conf/iclr/DosovitskiyB0WZ21], natural language processing [DBLP:conf/naacl/DevlinCLT19, DBLP:conf/nips/BrownMRSKDNSSAA20] and game playing [DBLP:journals/nature/SilverHMGSDSAPL16]. Given the remarkable success, DNNs are further deployed to safety-critical applications such as authentication [DBLP:journals/ijon/WangD21a] and autonomous driving [DBLP:journals/corr/BojarskiTDFFGJM16]. However, DNNs are proven to be vulnerable to a variety type of adversarial attacks. One notorious attack is the adversarial example [DBLP:journals/corr/GoodfellowSS14] that occurs at test-time, where an adversary fools well-trained DNN models by adding imperceptible perturbations on test images. Another well-known attack is the data poisoning attack [DBLP:conf/icml/BiggioNL12] that occurs in the training phase, where an adversary can inject well-crafted poisoned samples into the training data and introduce malicious behavior to models trained on the poisoned dataset. This paper focuses on backdoor attacks [gu2017badnets], where an adversary attempts to contaminate the training dataset via data poisoning and install a backdoor into models trained on the corrupted dataset. During inference, backdoored models
misclassify inputs with backdoor triggers to a target label while behaving normally on benign inputs.
@@
Backdoor attacks come in two flavors: dirty-label attacks [DBLP:journals/corr/abs-1712-05526, DBLP:conf/cvpr/Moosavi-Dezfooli17, DBLP:conf/iclr/NguyenT21] and clean-label attacks [DBLP:conf/icip/BarniKT19, DBLP:journals/corr/abs-1912-02771, DBLP:conf/aaai/SahaSP20, DBLP:conf/cvpr/HanZCL0L25]. Poisoned samples of dirty-label attacks are always mislabeled. For example, as shown in Figure 1, a poisoned sample “airplane” is labeled as “dog”. In contrast, clean-label attacks hold consistent labels to images content. Although dirty-label attacks are effective, they can be easily distinguished due to incorrect labels. Existing defenses have shown decent performance in detecting and mitigating such attacks [DBLP:journals/corr/abs-2012-10544, DBLP:journals/corr/abs-2007-08745]. On the other hand, clean-label attacks are more stealthy and insidious. The clean-label poisoned samples are almost visually indistinguishable from benign samples; thus, defenses against the attack become a more challenging task. Some recent works [DBLP:journals/corr/abs-2110-11571, DBLP:journals/corr/abs-2202-03423] attempted to alleviate the clean-label attacks by decoupling the training phase to suppress the backdoor injection. However, they require complicated algorithms with significantly more operations in training and do not identify the backdoor samples in the poisoned dataset. Thus, users have to re-run the defense algorithms every time they train a new model by using the same potentially poisoned dataset, which dramatically increases the cost.
@@
In this work, we propose UltraClean, a poisons-filtering-based framework (i.e., dataset cleanse) to detect backdoor samples, cleanse poisoned datasets, and train backdoor-mitigated models against both dirty-label and clean-label backdoor attacks. We focus on the image classification task. Our idea is inspired by prior works [DBLP:conf/iclr/LinGH19] and [DBLP:conf/cvpr/XieWMYH19, DBLP:conf/ndss/Xu0Q18], which have demonstrated that adversarial perturbations of adversarial examples are amplified during the feed-forward propagation (i.e., error amplification effect) in DNNs and can be effectively eliminated by simple image-denoising techniques. We argue that although backdoor samples hold fundamentally different generation mechanisms to adversarial examples, the backdoor triggers share a similar error amplification effect as the adversarial perturbations. To this end, our training framework, UltraClean, employs off-the-shelf image-denoising functions and the error implication effect to filter out poisoned samples from benign training data and thwart the backdoor implantation.
The proposed method first trains an arbitrary model on the potentially poisoned dataset and uses it as the backdoor detection model. We then produce two variants of each training image using denoising functions and feed the difference corresponding to the original image into the detection model. We leverage the error amplification effect to compute the susceptibility of the training data in a feed-forward pass. The susceptibility of poisoned data tends to be higher, guiding the model to detect and remove these samples. Finally, we obtain a backdoor-mitigated model by retraining on the sterilized dataset. We show that UltraClean is highly effective in defending against various representative attacks including both dirty-label and clean-label ones, achieving a high detection rate while maintaining the model accuracy.
@@
  • UltraClean is a once-for-all backdoor-free training framework and a poisons-filtering-based (i.e., dataset cleanse) defense against both dirty and clean-label backdoor attacks. The method not only defends against backdoor attacks but also identifies poisoned samples regardless of generation mechanisms and dataset complexity.
@@
  • We demonstrate that two widely adopted poisons-filtering-based defenses are only effective on dirty-label attacks but are unsuccessful in defending against clean-label attacks.
@@ 2.1 Backdoor Attacks on DNN
Backdoor attacks can be broadly categorized into data poisoning [gu2017badnets, DBLP:journals/corr/abs-1712-05526, DBLP:conf/aaai/SahaSP20, zhao2022towards, zhao2022clpa] and model poisoning [DBLP:conf/nips/NguyenT20, DBLP:conf/iclr/NguyenT21, DBLP:conf/iccv/DoanL0L21, DBLP:conf/nips/DoanLL21]. Data poisoning backdoor attacks implant the backdoors by modifying training data (i.e., injecting poisoned samples), which does not require access to the training process. Such attacks affect victims indirectly through the poisoned data set. In contrast, model poisoning backdoor attacks require full control of the training process to alter model parameters and embed backdoors into the model. Such attacks affect victims by providing the poisoned model directly. Our work focuses on the defense against data poisoning attacks via dataset cleanse.
The very first backdoor attack via data poisoning against neural networks is BadNets [gu2017badnets] where poisoned training samples are generated by stamping a pre-defined pattern (trigger) onto benign images and assigned with incorrect target labels (dirty-label). One drawback of BadNets is that the poisoned samples are suspicious and quite easy to detect upon human inspection. Later on, a series of works are proposed to enhance the stealthiness by improving the invisibility of backdoor triggers. [DBLP:journals/corr/abs-1712-05526] blends backdoor triggers with benign images; [DBLP:conf/codaspy/ZhongLSZ020] imposes imperceptible perturbations onto poisoned samples using universal adversarial perturbations [DBLP:conf/cvpr/Moosavi-Dezfooli17]; [DBLP:journals/tdsc/LiXZZZ21, li2021invisible] creates invisible triggers via steganography and regularization; [DBLP:conf/aaai/0005LMZ21] employs controlled detoxification to perform the attack in the feature space. However, all these methods are still dirty-label attacks and could be detected by inspecting data labels.
On the other hand, clean-label backdoor attacks are more stealthy and hard to detect since they retain consistent labels to the image contents. For instance, [DBLP:conf/icip/BarniKT19, DBLP:conf/eccv/LiuM0020, DBLP:conf/sp/QuiringR20] superimpose the trigger with benign images and employ different techniques (i.e., sinusoidal signal, reflection and image-scaling) to conceal the trigger; [DBLP:journals/corr/abs-1912-02771] creates poisoned sample by placing a stealthy patch on hard-to-classify images generated by generative models or adversarial perturbations; [DBLP:conf/aaai/SahaSP20] produces poisoned samples by minimizing their distance from source benign images in feature space. The objective of UltraClean is to defend against both dirty-label attacks and clean-label attacks in a simple and effective fashion.
@@ 2.2 Defenses
Defenses against backdoor attacks are extensively studied in recent years, which can be broadly categorized as training-phase defense and inference-phase defense. Training-phase defense attempts to eliminate backdoor before model deployment by filtering poisoned training samples [DBLP:conf/aaai/ChenCBLELMS19, DBLP:conf/nips/Tran0M18, DBLP:conf/sp/ChouTP20, DBLP:conf/uss/Tang0TZ21, DBLP:conf/acsac/GaoXW0RN19, DBLP:conf/iccv/ZengPMJ21, DBLP:journals/corr/abs-2104-11315], suppressing the effectiveness of poisoned data [DBLP:conf/iclr/DuJS20, DBLP:journals/corr/abs-2002-11497, DBLP:journals/corr/abs-2110-11571, DBLP:journals/corr/abs-2202-03423], or reconstructing the trained model [zhao2018resilience, DBLP:conf/raid/0017DG18, DBLP:conf/iclr/ZhaoCDRL20, DBLP:conf/iclr/LiLKLLM21]. Inference-phase defense alleviates the backdoor effect after model deployment by pre-processing data [DBLP:conf/acsac/DoanAR20, udeshi2019model, qiu2021deepsweep], filtering suspicious samples [javaheripi2020cleann, jin2020unified], synthesizing possible triggers [wang2019neural, chen2019deepinspect, DBLP:conf/nips/QiaoYL19, DBLP:conf/icdm/GuoWXXDS20, DBLP:conf/mm/ZhuNWXW20], or identifying if a model is backdoored [DBLP:journals/corr/abs-1911-07399, DBLP:conf/sp/XuWLBGL21, DBLP:conf/eccv/HuangPJT20, DBLP:conf/eccv/WangZLCXW20, DBLP:conf/cvpr/KolouriSPH20].
This paper proposes the first poisons-filtering-based defensive solution against a wide range of backdoor attacks (i.e., both dirty-label and clean-label). Our work is a training-phase defense where users seek to detect and remove malicious training samples and train a backdoor-free model. The most relevant works are the spectral signatures defense [DBLP:conf/nips/Tran0M18] and strong intentional perturbation defense [DBLP:conf/acsac/GaoXW0RN19] that can effectively distinguish poisoned samples in the dataset. However, we find that these defenses are only effective against dirty-label attacks.
@@ 3.1 Threat Model
We consider the same threat model defined in prior data poisoning backdoor works [DBLP:conf/nips/Tran0M18] where the adversary targets to implant a backdoor into a DNN for an image classification task by injecting a fraction of poisoned samples into the training dataset. We assume a strong adversary that knows standard training algorithms, classic model architectures, and the statistical information of the training dataset to craft powerful and stealthy poisoned samples. However, the model is not trained by the adversary but the user with possibly contaminated training dataset from an uncertified source. Under the threat model, we evaluate the effectiveness of the proposed defensive framework from three dimensions.
@@
Input: Training dataset $(\mathbf{x},\mathbf{y})$ $\sim$ $\boldsymbol{D_{tr}}$,
   Randomly initialized deep neural network model $\mathbf{M_{\theta}}$,
   Detection and removal threshold $\boldsymbol{{\beta}}$,
   Output of a potentially backdoored model $\boldsymbol{f_{\theta^{*}}(\cdot)}$,
   Total number of training samples $\boldsymbol{n}$
, Output: Sanitized training dataset $\boldsymbol{D_{s}}$,
    Post-clean model $\mathbf{M_{\hat{\theta}}}$
1
2#Train the model on the potentially poisoned dataset
3Minimize $\mathcal{L}=\mathds{E}_{(\mathbf{x},\mathbf{y})\sim\boldsymbol{D_{tr}}}[\ell(\mathbf{M_{\theta}}(\mathbf{x}),\mathbf{y})]\rightarrow\mathbf{M_{\theta^{*}}}$
4#Initialize the sanitized training set and score list
@@
6 $S$ $\leftarrow$ MaxHeap []
7#Generate denoised variants
8for all $\mathbf{x}$ do
9for $c_{i}$ in $\mathbf{x}$ do
10  $\tilde{{c_{i}}}$ $\frac{1}{C(p)}\sum_{\begin{subarray}{c}q\in\Omega(p,r)\end{subarray}}c_{i}(q)w(p,q)\rightarrow\tilde{\mathbf{x_{1}}}$
11  $\tilde{c_{i}}=median\{q\in\Omega(p):c_{i}(q)\}\rightarrow\tilde{\mathbf{x_{2}}}$
12end for
13#Enlarge error using error amplication effect
[…]
15#Compute susceptibility
[…]
18end for
19#Remove poisoned samples from the dataset
20for i = 0 to $\boldsymbol{{\beta}}\cdot\boldsymbol{n}$ do
[…]
22end for
23#Train/Retrain the model on the sterilized dataset
24Minimize $\mathcal{L}=\mathds{E}_{(\mathbf{x},\mathbf{y})\sim\boldsymbol{D_{s}}}[\ell(\mathbf{M_{\theta}}(\mathbf{x}),\mathbf{y})]\rightarrow\mathbf{M_{\hat{\theta}}}$
Return $\mathbf{M_{\hat{\theta}}},\boldsymbol{D_{s}}$.
@@ 11  $\tilde{c_{i}}=median\{q\in\Omega(p):c_{i}(q)\}\rightarrow\tilde{\mathbf{x_{2}}}$
Figure 2: Spatial and frequency domain views of two backdoor attacks. Top two rows are the dirty-label attack (BadNets [gu2017badnets]); Bottom two rows are the clean-label attack (Hidden Trigger Backdoor [DBLP:conf/aaai/SahaSP20]). As shown in the 4th and 8th columns, poisons reveal substantial qualitative noise difference than the clean counterparts. The noise difference in pixel space are amplified during the feed-forward propagation in a deep neural network and become a strong indicator to differentiate poisoned and benign samples.
@@ 3.3 Methodology
The key of UltraClean is to develop an approach that can differentiate poisoned and benign samples, which hence can be effective under both dirty-label and clean-label settings. In general, backdoor attacks require a trigger to induce DNNs to learn the malicious behavior. Such a trigger, whether visually perceptible or not, introduces extra noise (pixels that do not match the image content) to benign images. We plot the frequency-domain view in Figure 2 to visualize the signal distribution in images. The central area of the frequency-domain figure represents the low-frequency signal, while the peripheral area represents the high-frequency signal. As shown in the 1st and 5th columns, poisons demonstrate more high-frequency noise than benign data. The noise is bounded by the maximum perturbation added to the images and is sometimes hard to detect in the pixel space (i.e., some works attempt to minimize the perturbation [DBLP:journals/corr/abs-1912-02771, DBLP:conf/aaai/SahaSP20]). For example, it is barely impossible to perceptually distinguish the poison sample (the right side of the 3rd row) from the clean sample (the left side of the 3rd row) while they have obvious differences in the frequency domain (the 4th and 8th columns in the 4th row). Moreover, the noise distributions vary due to different mechanisms of poisoned sample generation, which demands high generalizability of the detection.
@@
Non-local mean denoising [DBLP:conf/cvpr/BuadesCM05] removes noise by replacing each pixel value with a weighted mean computed over global spatial regions. It is defined as:
$\tilde{c}(p)=\frac{1}{C(p)}\int g(d(\Omega(p,r),\Omega(q,r)))c(q)dq,$ (1)
where $\tilde{c}(p)$ is the denoised value of pixel $p$, $d(\cdot)$ represents the Euclidean distance between spatial regions $\Omega(p,r)$ and $\Omega(q,r)$. $\Omega(p,r)$ and $\Omega(q,r)$ are search windows centered at pixels $p$ and $q$, and the boundary of which is defined by $r$. $C(\cdot)$ is a normalization function and $g(\cdot)$ is a decreasing function. In this work, we exploit the pixel-wise implementation of the algorithm. Consider a color image with RGB channels as $\mathbf{x}=(c_{1},c_{2},c_{3})$, the denoising operation is expressed as:
@@
We acknowledge there exist many other denoising algorithms designed specifically for certain types of noise and conduct an abliation study in Appendix G, where we observe that using the aforementioned two denoising techniques can already achieve excellent performance with trivial complexity. The workflow of UltraClean is straightforward, for each training image $\mathbf{x}$, we generate two denoised versions $\tilde{\mathbf{x_{1}}}$ (e.g., the 2nd and 6th columns in Fig 2) and $\tilde{\mathbf{x_{2}}}$ (e.g., the 3rd and 7th columns). $\tilde{\mathbf{x_{1}}}$ and $\tilde{\mathbf{x_{2}}}$ serve as baselines to compute the susceptibility. We define susceptibility as the error (i.e., noise difference) between original images and denoised variants. It can be seen from the 4th and 8th columns that the error of poisoned images are considerably higher than that of clean images. However, as seen in the 6th and 7th columns, the denoising functions do not significantly reduce the noise in the denoised variants compared to the original poisoned samples (the 5th column), so we are not able to distinguish poisoned samples from benign samples by simply computing the susceptibility in the pixel space.
We propose a method to enlarge the error to facilitate the computation using the error amplification effect, which is a property of DNNs that minor adversarial perturbations in model inputs accumulate during forward propagation and affect the model outputs [DBLP:conf/iclr/LinGH19]. We feed the original image and denoised variants into the pre-trained DNN and obtain three vectors $\mathbf{v}$, $\mathbf{v_{1}}$, and $\mathbf{v_{2}}$, which are the softmax outputs of the model. We compute the $\ell_{1}$-norm distances of ($\mathbf{v}$, $\mathbf{v_{1}}$) and ($\mathbf{v}$, $\mathbf{v_{2}}$), respectively, and then aggregate the results to obtain the amplified noise difference. The susceptibility can thus be mathematically defined as:
[…]